OpenStudy (anonymous):

Given two rectangles, one with length = l and width = w, and the second rectangle with length = 2l and width = 4w. How much larger will the perimeter of the second rectangle be than the perimeter of the first rectangle?

OpenStudy (anonymous):

A. P2 - P1 = 2 * (l + 3w) B. P2 = 2P1 + w C. P2 = 2P1 + 4w D. P2 = P1 + w

OpenStudy (rose21):

C. p2=2p1+4w